# Mining Your Own Secrets: Diffusion Classifier Scores for Continual Personalization of Text-to-Image Diffusion Models

### v1

@@ 4 Experiments
Implementation. We use the Stable Diffusion v1.4 (Rombach et al., 2022) as our backbone based on the Diffusers library (von Platen et al., 2022). Following CD (Kumari et al., 2023), we train all the models for $1000$ iterations on all but the Celeb-A setup, where we use 2000 training iterations to capture more fine-grained facial attributes (Smith et al., 2024b). In favor of zero-shot generalization, we finetune our hyperparameters only on the six task sequence of Custom Concept. For both EWC and DSC, we set the number of consolidation iterations to ${1/5}^{\text{th}}$ of that of the training iterations number. For computing DC scores, the temperature $\tau$ is set to 1.0 for all but the teacher LoRA in DSC where we set $\tau$ to 0.05. The cardinality of $\mathbf{c}^{k}$ for EWC is set to 3. The DSC loss weights $\gamma$ and $\lambda$ are set to 0.1 and 1.5, respectively. We detail on implementation and hyperparameters in App. H.
Evaluation. For each concept, we use DDPM sampling with 50 inference steps to generate 400 images using the prompt “a photo of a $\text{V}^{*}_{i}$”, where $\text{V}_{i}^{*}$ is the modifier token learned for the $i^{\text{th}}$ task (Kumari et al., 2023). For CD and TI, we additionally include the concept name after the modifier token. We encode the generated and the target (real) images using CLIP image encoder (Radford et al., 2021a), and use the features for computing our metrics. We quantify a task’s performance using the following metrics computed as the average over all seen concepts: (a) $A_{\text{MMD}}$: the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) between real and target image feature distribution, (b) CLIP I2I: the CLIP image-alignment (Gal et al., 2023) and (c) KID: the Kernel Inception Distance (KID) (Bińkowski et al., 2018). To quantify forgetting, we use the forgetting metric $F_{\text{MMD}}$ (Smith et al., 2024b), which is the average change in past task concept generations from task to task. Given the relative nature of $F_{\text{MMD}}$, we opt for a more robust backward transfer of MMD: $\text{BwT}_{\text{MMD}}$, which measures the absolute change in MMD scores for previous concept generations. We also mention the percentage of parameters trained for a task as $\mathsf{N_{\text{param}}\;Train}$, and the percentage of parameters stored between tasks as $\mathsf{N_{\text{param}}\;Store}$. We detail these metrics in App. E.
@@ 4.1 Results
Continual Personalization Of Custom Concepts. We evaluate our method on the Custom Concept dataset (Kumari et al., 2023) that comprises diverse categories such as plushies, wearables, and toys (see App. D). We randomly sample six such concepts to form six tasks. Figure 4(a) shows the samples generated for tasks 1, 3, and 6 after training on all 6 tasks: the upper row compares the baselines while the lower row compares our variants. We report the quantitative results in Table 1. Marked by its high backward transfer scores, we find that CD (sequential) is prone to forgetting the category attributes such as the color, background, and appearance while also struggling with plasticity loss that leads to the generation of incorrect backgrounds for task 6. While Textual Inversion Gal et al. (2023) has zero forgetting (due to frozen backbone), it remains poor at capturing the custom category. LoRA (sequential) undergoes catastrophic forgetting of past tasks’ concepts (it has the highest BwT scores among LoRA-based methods) while C-LoRA displays a higher degree of interference for past tasks. Both these methods have reduced plasticity for acquiring the last task images – sunglasses appear in incorrect numbers/styles/backgrounds. LoRA with EWC shows an improvement on both plasticity and forgetting. Incorporating DC scores into it further helps reduce the forgetting, with an improvement of 3.19 points for $\text{BwT}_{\text{MMD}}$. On the contrary, a naive DSC undergoes huge forgetting as it performs distillation without the previous task data. Using DSC with EWC helps remedy this forgetting, albeit at the cost of reduced plasticity, e.g., task 6 sunglasses with three pairs of glasses. Incorporating DC scores with DSC EWC further helps improve the results, and is our best performing variant as it displays the least KID, $A_{\text{MMD}}$, and $\text{BwT}_{\text{MMD}}$ scores.
@@
Method $\mathsf{N_{\text{param}}}\;$ Train (↓) $\mathsf{N_{\text{param}}}\;$ Store (↓) KID (↓) (x $10^{3}$) $A_{\text{MMD}}$ (↓) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ (↓) CLIP I2I (↓) (x100) $F_{\text{MMD}}$ (↓)
@@
Method $\mathsf{N_{\text{param}}}\;$ Train (↓) $\mathsf{N_{\text{param}}}\;$ Store (↓) KID (↓) (x $10^{3}$) $A_{\text{MMD}}$ (↓) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ (↓) CLIP I2I (↓) (x100) $F_{\text{MMD}}$ (↓)
@@
Continual Personalization of Household objects and Real Faces. We further compare the LoRA-based methods on finegrained images of household objects from the Textual Inversion (TI) dataset (Gal et al., 2023), and that of celebrity faces from the CelebA dataset (Liu et al., 2015). These setups comprise task sequences of length 9 and 10, respectively (see App. D for details). We leave the qualitative and quantitative results in App. F. We find the results to follow the same pattern as with Custom Concept and Landmarks. For both these setups, C-LoRA fails to remember the finegrained details of previous concepts, e.g., exact object/facial attributes, and instead, only retains the overall feature, e.g., the dominant color. Also, our DSC EWC DC variant performs the best here.
@@ 4.2 Long task sequence
We study the scalability of our proposed methods to a sequence of 50 concepts chosen at random from the Custom Concept dataset (Kumari et al., 2023), with variable number of training images per concept. To avoid any learning bias from large early tasks with more training images, we pick all 50 tasks at random rather than adding them over our six tasks sequence. Fig. 5 compares the results of C-LoRA, LoRA EWC, and Ours (EWC DC). Similar to Smith et al. (2024b), we find that the performance of C-LoRA saturates as the number of tasks grow. Instead, applying EWC on the LoRA parameters emerges as a better performer on the long run. EWC with DC retains the performance particularly on the latter ($>35$) tasks (see App. fig. 15 for qualitative comparison).
@@ F.2 Results on Household objects and Real Faces
Table 3: TI results after 9 tasks (avg. over 3 seeds): (↓) indicates lower is better.
Method $\mathsf{N_{\text{param}}}\;$ Train (↓) $\mathsf{N_{\text{param}}}\;$ Store (↓) KID (↓) (x $10^{3}$) $A_{\text{MMD}}$ (↓) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ (↓) CLIP I2I (↓) (x100) $F_{\text{MMD}}$ (↓)
CD 2.23 100.0 68.13 296.11 187.4 0.025 -204.31
LoRA sequential 0.09 100.0 67.1 261.07 132.58 0.008 -125.03
C-LoRA 0.09 100.81 69.04 181.33 105.7 0.047 -106.16
EWC 0.09 100.81 71.46 158.93 88.21 0.003 -93.18
Ours: EWC DC 0.09 100.81 73.08 139.66 82.75 0.047 -79.0
Ours: DSC 0.09 100.81 69.81 168.09 96.41 0.008 -135.74
Ours: DSC EWC 0.09 100.81 71.08 135.61 80.6 0.008 -84.11
Ours: DSC EWC DC 0.09 100.81 72. 66 121.08 78.34 0.005 -73.18
Table 4: Celeb-A results after 10 tasks (avg. over 3 seeds): (↓) indicates lower is better.
Method $\mathsf{N_{\text{param}}}\;$ Train (↓) $\mathsf{N_{\text{param}}}\;$ Store (↓) KID (↓) (x $10^{3}$) $A_{\text{MMD}}$ (↓) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ (↓) CLIP I2I (↓) (x100) $F_{\text{MMD}}$ (↓)
@@ G.1 Compatibility with VeRA
Method $\mathsf{N_{\text{param}}}\;$ Train (↓) $\mathsf{N_{\text{param}}}\;$ Store (↓) KID (↓) (x $10^{3}$) $A_{\text{MMD}}$ (↓) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ (↓) CLIP I2I (↓) (x100) $F_{\text{MMD}}$ (↓)
@@ H.1 Implementation Details
We use the Hugging Face Accelerate library (Gugger et al., 2022) for distributed training/inference of our models. Our experiments are conducted using four RTX A6000 GPUs with 48 GB memory each. For all the compared methods, we set the batch size to 1 during training and inference. To allow for larger effective batch sizes during training, we set the gradient accumulation steps to 8. For a fair comparison with CD, we perform regularization during training using an auxiliary dataset of $200$ images generated by the pretrained backbone using the prompt “a photo of a person” (Smith et al., 2024b). We follow C-LoRA for preserving a number of hyperparameters setup. Namely, we set the LoRA rank to 16, and the EWC loss coefficient to $61e$ 6 for all our experiments. We use a learning rate of $65e-$ 6 all non-LoRA methods, and a learning rate of $45e-$ 4 for all LoRA-based methods. We implemented C-LoRA from scratch and following the authors, used a coefficient of $81e$ 8 for the self-regularization loss $\mathcal{L}_{\text{forget}}$ (see Eq. 2). Lastly, following the default setup of the Diffusers library (von Platen et al., 2022), we set the classifier guidance scale to 7.5 to allow for a higher adherence to the conditional signal.
@@ Appendix I Training Time Complexity Analysis
We use the soft-O notation $\mathcal{\tilde{O}}$ (Van Rooij et al., 2019) to describe the time complexity while ignoring the logarithmic factors. Formally, for some constant $k$, $\mathcal{\tilde{O}}(f(n))=\mathcal{O}\big{(}f(n)*\log^{k(n)}\big{)}$ provides the upper bound for $f$, like the standard big-O notation $\mathcal{O}$ but hides the factors involving powers of logarithms, i.e., $\mathcal{\tilde{O}}(n)$ could represent $\mathcal{O}(n\log n),\mathcal{O}(n\log\log n),\mathcal{O}(n\log^{2}n)$, etc.
As also stated in the main paper, we consider a continual personalization setup with $N$ number of tasks such that each task comprises on new concept to acquire. For the ease of computation, we assume that each task has a fixed number of training images, $|\mathcal{D}|$. Note that for our CL setup, we use the same number of training epochs for each task. This lets us ignore the factor of training epochs in deriving the training time complexity. Lastly, since both C-LoRA (Smith et al., 2024b) and our setup train a single LoRA and a modifier token per task, their complexity of a forward pass remains the same, and can be safely ignored. Put together, we can state time complexity as a function that grows linearly with more training samples $|\mathcal{D}|$. We list the training time complexities of the compared methods in Table 6 and detail their derivation below:

### v2

@@ 4 Experiments
Implementation. We use the Stable Diffusion v1.4 (Rombach et al., 2022) as our backbone based on the Diffusers library (von Platen et al., 2022). Following CD (Kumari et al., 2023), we train all the models for $1000$ iterations on all but the Celeb-A setup, where we use 2000 training iterations to capture more fine-grained facial attributes (Smith et al., 2024b). In favor of zero-shot generalization, we finetune our hyperparameters only on the six task sequence of Custom Concept. For both EWC and DSC, we set the number of consolidation iterations to ${1/5}^{\text{th}}$ of that of the training iterations number. For computing DC scores, the temperature $\tau$ is set to 1.0 for all but the teacher LoRA in DSC where we set $\tau$ to 0.05. The cardinality of $\mathbf{c}^{k}$ for EWC is set to 5. The DSC loss weights $\gamma$ and $\lambda$ are set to 0.1 and 1.5, respectively. We detail on implementation and hyperparameters in App. H.
Evaluation. For each concept, we use DDPM sampling with 50 inference steps to generate 400 images using the prompt “a photo of a $\text{V}^{*}_{i}$”, where $\text{V}_{i}^{*}$ is the modifier token learned for the $i^{\text{th}}$ task (Kumari et al., 2023). For CD and TI, we additionally include the concept name after the modifier token. We encode the generated and the target (real) images using CLIP image encoder (Radford et al., 2021a), and use the features for computing our metrics. We quantify a task’s performance using the following metrics computed as the average over all seen concepts: (a) $A_{\text{MMD}}$: the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) between target and synthetic image feature distribution, (b) CLIP I2I: the CLIP image-alignment (Gal et al., 2023) and (c) KID: the Kernel Inception Distance (KID) (Bińkowski et al., 2018). To quantify forgetting, we use the forgetting metric $F_{\text{MMD}}$ (Smith et al., 2024b), which is the average change in past task concept generations from task to task. Given the relative nature of $F_{\text{MMD}}$, we opt for a more robust backward transfer of MMD: $\text{BwT}_{\text{MMD}}$, which measures the absolute change in MMD scores for previous concept generations. We also mention the percentage (wrt the U-Net) of parameters trained for a task as $\mathsf{N_{\text{param}}\;Train}$, and the percentage of parameters stored between tasks as $\mathsf{N_{\text{param}}\;Store}$. We detail these metrics in App. E.
@@ 4.1 Results
Continual Personalization of Custom Concepts. We evaluate our method on the Custom Concept dataset (Kumari et al., 2023) that comprises diverse categories such as plushies, wearables, and toys (see App. D). We randomly sample six such concepts to form six tasks. Figure 4(a) shows the samples generated for tasks 1, 3, and 6 after training on all 6 tasks: the upper row compares the baselines while the lower row compares our variants. We report the quantitative results in Table 1. Marked by its low $\text{BwT}_{\text{MMD}}$ scores, we find that CD is prone to forgetting the category attributes such as the color, background, and appearance while also struggling with plasticity loss that leads to the generation of incorrect backgrounds for task 6. While Textual Inversion Gal et al. (2023) has zero forgetting due to its frozen backbone, it remains poor at acquiring categories. LoRA (sequential) undergoes catastrophic forgetting of past tasks’ concepts (it has the highest $\text{BwT}_{\text{MMD}}$ scores among LoRA-based methods) while C-LoRA displays a higher degree of interference for past tasks. Both these methods have reduced plasticity for acquiring the last task images – sunglasses appear in incorrect numbers/styles. LoRA with EWC shows an improvement on both plasticity and forgetting. Incorporating DC scores into it further helps reduce the forgetting, with an improvement of 3.19 points for $\text{BwT}_{\text{MMD}}$. On the contrary, a naive DSC undergoes huge forgetting as it performs distillation without the previous task data. Using DSC with EWC helps remedy this forgetting, albeit at the cost of reduced plasticity, e.g., task 6 sunglasses with three pairs of glasses. Incorporating DC scores with DSC EWC further helps improve on these results, and is our best performing variant as it displays the least KID, $A_{\text{MMD}}$, and $\text{BwT}_{\text{MMD}}$ scores (see App. F for taskwise performances).
@@
Method $\mathsf{N_{\text{param}}}\;$ $\mathsf{Train}$($\downarrow$) $\mathsf{N_{\text{param}}}\;$ $\mathsf{Store}$ ($\downarrow$) KID ($\downarrow$) (x $10^{5}$) $A_{\text{MMD}}$ ($\downarrow$) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ ($\uparrow$) CLIP I2I ($\uparrow$) (x100) $F_{\text{MMD}}$ ($\downarrow$)
@@
Method $\mathsf{N_{\text{param}}}\;$ $\mathsf{Train}$($\downarrow$) $\mathsf{N_{\text{param}}}\;$ $\mathsf{Store}$ ($\downarrow$) KID ($\downarrow$) (x $10^{5}$) $A_{\text{MMD}}$ ($\downarrow$) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ ($\uparrow$) CLIP I2I ($\uparrow$) (x100) $F_{\text{MMD}}$ ($\downarrow$)
@@
Continual Personalization of Household objects and Real Faces. We further compare the LoRA-based methods on finegrained images of household objects from the Textual Inversion (TI) dataset (Gal et al., 2023), and that of celebrity faces from the CelebA dataset (Liu et al., 2015). These setups comprise task sequences of length 9 and 10, respectively (see App. D for details). We leave the qualitative and quantitative results in App. F.2. We find the results to follow the same pattern as with Custom Concept and Landmarks. For both these setups, C-LoRA fails to remember the finegrained details of previous concepts, e.g., exact object/facial attributes, and instead, only retains the overall feature, e.g., the dominant color. Also, our DSC EWC DC variant performs the best here.
@@ 4.2 Long task sequence
We study the scalability of our proposed methods to a sequence of 50 concepts chosen at random from the Custom Concept dataset (Kumari et al., 2023), with variable number of training images per concept. To avoid any learning bias from large early tasks with more training images, we pick all 50 tasks at random rather than adding them over our six tasks sequence. Fig. 5 compares the results of C-LoRA, LoRA EWC, and Ours (EWC DC). Similar to Smith et al. (2024b), we find that the performance of C-LoRA saturates as the number of tasks grow. Instead, applying EWC on the LoRA parameters emerges as a better performer on the long run. EWC with DC retains the performance particularly on the latter ($>35$) tasks (see App. F.1 for qualitative comparison).
@@ Appendix E Metrics Definition
In terms of magnitude, the higher the better ($\uparrow$) holds for: the image alignment (I2I) and the backward transfer ($\text{BwT}_{\text{MMD}}$) metrics, while the lower the better ($\downarrow$) holds for all other metrics.
@@ F.2 Results on Household objects and Real Faces
Table 3: TI results after 9 tasks (avg. over 3 seeds): ($\downarrow$) indicates lower is better.
Method $\mathsf{N_{\text{param}}}\;$ $\mathsf{Train}$($\downarrow$) $\mathsf{N_{\text{param}}}\;$ $\mathsf{Store}$ ($\downarrow$) KID ($\downarrow$) (x $10^{5}$) $A_{\text{MMD}}$ ($\downarrow$) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ ($\uparrow$) CLIP I2I ($\uparrow$) (x100) $F_{\text{MMD}}$ ($\downarrow$)
CD 2.23 100.0 296.11 187.4 -204.31 68.13 0.025
LoRA sequential 0.09 100.0 261.07 132.58 -125.03 67.1 0.008
C-LoRA 0.09 100.81 181.33 105.7 -106.16 69.04 0.047
EWC 0.09 100.81 158.93 88.21 -93.18 71.46 0.003
Ours: EWC DC 0.09 100.81 139.66 82.75 -79.0 73.08 0.047
Ours: DSC 0.09 100.81 168.09 96.41 -135.74 69.81 0.008
Ours: DSC EWC 0.09 100.81 135.61 80.6 -84.11 71.08 0.008
Ours: DSC EWC DC 0.09 100.81 121.08 78.34 -73.18 72. 66 0.005
Table 4: Celeb-A results after 10 tasks (avg. over 3 seeds): ($\downarrow$) indicates lower is better.
Method $\mathsf{N_{\text{param}}}\;$ $\mathsf{Train}$($\downarrow$) $\mathsf{N_{\text{param}}}\;$ $\mathsf{Store}$ ($\downarrow$) KID ($\downarrow$) (x $10^{5}$) $A_{\text{MMD}}$ ($\downarrow$) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ ($\uparrow$) CLIP I2I ($\uparrow$) (x100) $F_{\text{MMD}}$ ($\downarrow$)
@@ G.1 Compatibility with VeRA
Method $\mathsf{N_{\text{param}}}\;$ $\mathsf{Train}$($\downarrow$) $\mathsf{N_{\text{param}}}\;$ $\mathsf{Store}$ ($\downarrow$) KID ($\downarrow$) (x $10^{5}$) $A_{\text{MMD}}$ ($\downarrow$) (x $10^{3}$) $\text{BwT}_{\text{MMD}}$ ($\uparrow$) CLIP I2I ($\uparrow$) (x100) $F_{\text{MMD}}$ ($\downarrow$)
@@ H.1 Implementation Details
We use the Hugging Face Accelerate library (Gugger et al., 2022) for distributed training/inference of our models. Our experiments are conducted using four RTX A6000 GPUs with 48 GB memory each. For all the compared methods, we set the batch size to 1 during training and inference. To allow for larger effective batch sizes during training, we set the gradient accumulation steps to 8. For a fair comparison with CD, we perform regularization during training using an auxiliary dataset of $200$ images generated by the pretrained backbone using the prompt “a photo of a person” (Smith et al., 2024b). We follow C-LoRA for preserving a number of hyperparameters setup. Namely, we set the LoRA rank to 16, and the EWC loss coefficient to $61e$ 6 for all our experiments. We use a learning rate of $65e-$ 6 for all non-LoRA methods, and a learning rate of $45e-$ 4 for all LoRA-based methods. We implemented C-LoRA from scratch and following the authors, used a coefficient of $81e$ 8 for the self-regularization loss $\mathcal{L}_{\text{forget}}$ (see Eq. 2). Lastly, following the default setup of the Diffusers library (von Platen et al., 2022), we set the classifier guidance scale to 7.5 to allow for a higher adherence to the conditional signal.
@@ Appendix I Training Time Complexity Analysis
We use the soft-O notation $\mathcal{\tilde{O}}$ (Van Rooij et al., 2019) to describe the time complexity while safely ignoring the logarithmic factors. Formally, for some constant $k$, $\mathcal{\tilde{O}}(f(n))=\mathcal{O}\big{(}f(n)*\log^{k(n)}\big{)}$ provides the upper bound for $f$, like the standard big-O notation $\mathcal{O}$ but hides the factors involving powers of logarithms, i.e., $\mathcal{\tilde{O}}(n)$ could represent $\mathcal{O}(n\log n),\mathcal{O}(n\log\log n),\mathcal{O}(n\log^{2}n)$, etc.
As also stated in the main paper, we consider a continual personalization setup with $N$ number of tasks such that each task comprises on new concept to acquire. For the ease of computation, we assume that each task has a fixed number of training images, $|\mathcal{D}|$. Note that for our CL setup, we use the same number of training epochs for each task. This lets us ignore the factor of training epochs in deriving the training time complexity. Lastly, since both C-LoRA (Smith et al., 2024b) and our setup train a single LoRA and a modifier token per task, their complexity of a forward pass remains the same, and can be safely ignored. Put together, we can state time complexity as a function that grows linearly with more training samples $|\mathcal{D}|$. We list the training time complexities of the compared methods in Table 6 and detail on their derivation below: